OpenStudy (anonymous):

Simplify completely:

OpenStudy (anonymous):

\[8x^3+6x^2-4x \over -2x\]

jimthompson5910 (jim_thompson5910):

\[\Large \frac{8x^3+6x^2-4x}{-2x}\] \[\Large \frac{-2x(-4x^2-3x+2)}{-2x}\] \[\Large \frac{\cancel{-2x}(-4x^2-3x+2)}{\cancel{-2x}}\] \[\Large -4x^2-3x+2\] So \[\Large \frac{8x^3+6x^2-4x}{-2x}=-4x^2-3x+2\] for all values of x where x is not zero.

OpenStudy (anonymous):

thanks you are a lifesaver

jimthompson5910 (jim_thompson5910):

yw